OpenStudy (anonymous):

Let {v1, v2, v3} be a set of nonzero vectors in R^m such that tr(vi)vj = 0 when i does not equal j. Show that the set is linearly independent.

OpenStudy (anonymous):

\[v _{i}^{T}v_{j} = 0 \] when \[i \neq j\] Sorry if my notation for this stuff is bad, I'm new to it is all.

OpenStudy (helder_edwin):

let \(\alpha,\beta,\gamma\) such that \[\large \alpha v_1+\beta v_2+\gamma v_3=0. \] Then \[\large v_1^t(\alpha v_1+\beta v_2+\gamma v_3)=v_1^t0=0 \] \[\large v_1^t(\alpha v_1)+v_1^t(\beta v_2)+v_1^t(\gamma)v_3=0 \] \[\large \alpha(v_1^tv_1)+\beta(v_1^tv_2)+\gamma(v_1^tv_3)=0 \] since all are mutualy orthogonal \[\large \alpha(v_1^tv_1)=0 \] since \(v_1\neq0\) then \(\alpha=0\). The same can be dome for \(v_2\) and \(v_3\). So the \(v_i\) are linearly independent.